Use back-substitution to solve the system of linear equations. (If there is no solution, enter NO SOLUTION. If the system is dependent, set z = a and solve for x and y in terms of a.)

6x − 5y + z = 52
−y + z = 11
z = 6

(x, y, z) =
Answer #1

Solution given system of equations 6N-5y + 2 = 52 - Y + 2 = 11 2 = 6 2 3 using back. substitution. from equation ③ Z=6 y valu